OpenStudy (anonymous):

400 draws are made at random with replacement from 5 tickets that are marked -2, -1, 0, 1, and 2 respectively. Find the expected value of the sum of the positive numbers drawn

OpenStudy (kropot72):

The probability of drawing a ticket marked 1 on each draw is 1/5. The probability of drawing a ticket marked 2 on each draw is 1/5. The expected number of tickets marked 1 drawn in 400 draws is\[400\times \frac{1}{5}\] The expected number of tickets marked 2 drawn in 400 draws is \[400\times \frac{1}{5}\] The expected value of the sum of the positive numbers drawn is \[(1\times 400\times \frac{1}{5})+(2\times 400\times \frac{1}{5})=you\ can\ calculate\]

OpenStudy (anonymous):

240

OpenStudy (anonymous):

ok i have now seen where i was wrong. i just obtained the number of times a positive appears which=np=(2/5)×400=160. n later multiplied by the average of the positives. ie (1+2)/2. Thanks for showing me the way !!!

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

So The expected value of the sum of the all drawn is ? \[ \left( 1*400*\frac{ 1 }{ 5 } \right)+\left( 2*400*\frac{ 1 }{ 5 } \right)+\left( 3*400*\frac{ 1 }{ 5 } \right)+\left( 4*400*\frac{ 1 }{ 5 } \right)+\left( 5*400*\frac{ 1 }{ 5 } \right)\]

OpenStudy (anonymous):

all the numbers drawn